# Slow convergence of spin-wave expansion and magnon dispersion in the 1/3 plateau of the triangular XXZ antiferromagnet

### v1

@@ Abstract
The nearest-neighbour spin-$1/2$ XXZ antiferromagnet on the triangular lattice is one of the prototype models in the theory of geometrically-frustrated quantum magnetism, and has been a subject of investigation for decades. In the case in which the exchange anisotropy is of easy-axis type, an early analysis [1] suggested the possibility that the model may host a quantum spin liquid ground state, but later studies did not corroborate this scenario and predicted instead an ordered state at zero field, followed by a sequence of transitions between ordered phases in presence of an external magnetic field directed along the $c$ axis (the easy axis of the exchange anisotropy) [2, 3, 4, 5, 6, 7, 8, 9, 10]. For zero magnetic field, in particular, different theoretical analyses predicted a “spin-supersolid” phase [2, 3, 4, 5, 6, 7, 8, 9, 10], with simultaneous breaking of a $Z_{3}$ translational invariance and the $U(1)$ rotational symmetry. Some results reported recently, however, gave indications that the ground state may be a “solid” with unbroken $U(1)$ symmetry [11, 12] (see also Ref. [10]). In presence of a $c$-axis oriented field, the zero-temperature phase diagram has been predicted to present, in addition to the low-field phase, a $1/3$-plateau phase, a high-field supersolid, and a fully saturated phase [7, 9, 13, 10].
@@
Figure 1: $uud$ configuration and magnon hoppings in second-order perturbation theory. At order $\alpha^{2}$, the $u$ excitations are described by an effective tight binding problem with hopping amplitudes $t_{1u}$, $t_{2u}$, and $t_{3u}$ on the honeycomb lattice. The $d$ excitations, at the same order, are described by an effective hopping problem with amplitude $t_{1d}$ between the nearest-neighbour C sites (which are second-nearest neighbours on the full lattice).
@@
Turning on a small $J_{xy}$ leads to a perturbative dressing of the $uud$ ground state and induces a dispersion of the magnon excitations [21]. For the $u$ excitations, the dispersion at second order in $\alpha$ is described by an effective hopping problem on the honeycomb lattice, with three non-zero hopping amplitudes between first, second, and third nearest neighbours (see Fig. 1). By an explicit calculation, we find that the corresponding hopping amplitudes are $t_{1u}=-\alpha(1-\alpha/2)$, $t_{2u}=t_{3u}=\alpha^{2}J_{zz}/4$. In addition, perturbative corrections give a momentum-independent energy shift, which arises from the disturbance induced by an excitation on the virtual transitions occurring in its proximity. This leads to a renormalization of the gap, which can be encoded in an energy shift $V_{u}=-5\alpha^{2}J_{zz}/8$. We note that the effective parameters $t_{1u}$, $t_{2u}$, $t_{3u}$, $V_{u}$ which we derived are equal, in a different notation, to those obtained in Ref. [6] for an equivalent problem of hard-core bosons.
@@
$\begin{split}\epsilon^{(2,{\rm LSWT})}_{u,\pm}(\mathbf{k})&=g_{z}\mu_{\rm B}H-\frac{S}{6}\alpha^{2}J_{zz}|f_{\mathbf{k}}|^{2}\\ &\qquad\pm S\alpha J_{zz}\left|f_{\mathbf{k}}-\frac{\alpha}{6}f^{*2}_{\mathbf{k}}\right|~{},\\ \epsilon^{(2,{\rm LSWT})}_{d,\pm}(\mathbf{k})&=6SJ_{zz}-g_{z}\mu_{\rm B}H-\frac{S}{3}\alpha^{2}J_{zz}|f_{\mathbf{k}}|^{2}~{}.\end{split}$ (5)
@@
In the concluding page, we note that the expansion near the Ising limit offers the possibility to assess not only the quality of the LSWT approximation, but also the convergence of the semiclassical $1/S$ series. This can be controlled by studying the generalization of the problem to an arbitrary value of the spin $S$. To repeat the analysis, the $uud$ ground state configuration is promoted to a state in which two sublattices (A and B) are occupied by spins $+S$, whereas the remaining sublattice C is occupied by spins $-S$, and the $u$ and $d$ excitations, become the lowering or raising by $1$ of the spins on the A, B or on the C sublattices. At order $\alpha^{2}$, the effective hopping problem on the honeycomb lattice which governs the $u$ excitations has the same hopping amplitudes as in the $S=1/2$ case, but with a nontrivial dependence on the spin $S$:
@@
$\begin{split}\epsilon^{(2)}_{u,\pm}(\mathbf{k})&=g_{z}\mu_{\rm B}H+V_{u}+t_{1u}\Big{[}-\gamma(|f_{\mathbf{k}}|^{2}-3)\\ &\pm\left|(1-2\gamma)_{\mathbf{k}}+\gamma f^{*2}_{\mathbf{k}}\right|\Big{]}~{},\end{split}$ (15)
@@
After applying the $\pi$ rotation $S_{i}^{x}=S_{i}^{x}$, $S_{i}^{y}=-S^{y}_{i}$, $S^{z}_{i}=-S_{i}^{z}$ to the spins residing on the C sublattice, introducing the Holstein-Primakoff representations
@@
11footnotetext: See the Supplemental Materials of Ref. [20].

### v2

@@ Abstract
The nearest-neighbour spin-$1/2$ XXZ antiferromagnet on the triangular lattice is one of the prototype models in the theory of geometrically-frustrated quantum magnetism, and has been a subject of investigation for decades. In the case in which the exchange anisotropy is of easy-axis type, an early analysis [1] suggested the possibility that the model may host a quantum spin liquid ground state, but later studies did not corroborate this scenario and predicted instead an ordered state at zero field, followed by a sequence of transitions between ordered phases in presence of an external magnetic field directed along the $c$ axis (the easy axis of the exchange anisotropy) [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. For zero magnetic field, in particular, different theoretical analyses predicted a “spin-supersolid” phase [4, 5, 6, 7, 8, 9, 10, 11, 12], with simultaneous breaking of a $Z_{3}$ translational invariance and the $U(1)$ rotational symmetry. Some results reported recently, however, gave indications that the ground state may be a “solid” with unbroken $U(1)$ symmetry [13, 14] (see also Ref. [12]). In presence of a $c$-axis oriented field, the zero-temperature phase diagram has been predicted to present, in addition to the low-field phase, a $1/3$-plateau phase, a high-field supersolid, and a fully saturated phase [9, 11, 15, 12].
@@
Figure 1: (a) $uud$ configuration and magnon hoppings in second-order perturbation theory. At order $\alpha^{2}$, the $u$ excitations are described by an effective tight binding problem with hopping amplitudes $t_{1u}$, $t_{2u}$, and $t_{3u}$ on the honeycomb lattice. The $d$ excitations, at the same order, are described by an effective hopping problem with amplitude $t_{1d}$ between the nearest-neighbour C sites (which are second-nearest neighbours on the full lattice). (b) Perturbative process leading to the second-order correction to $t_{1u}$: two pairs of spins are interchanged sequentially (flipping pair ’1’, then pair ’2’) resulting in the motion of the magnon excitation. $t_{2u}$, $t_{3u}$, and $t_{1d}$ are determined by similar processes.
@@
Turning on a small $J_{xy}$ leads to a perturbative dressing of the $uud$ ground state and induces a dispersion of the magnon excitations [23]. For the $u$ excitations, the dispersion at second order in $\alpha$ is described by an effective hopping problem on the honeycomb lattice, with three non-zero hopping amplitudes between first, second, and third nearest neighbours (see Fig. 1). By an explicit calculation, we find that the correspond 22footnotetext: A. Mauri and F. Mila, ’Spin-wave expansion and magnon dispersion in the 1/3 plateau of the triangular XXZ antiferromagnet’. Zenodo, 2025. doi: https://doi.org/10.5281/zenodo.15270189ing hopping amplitudes are $t_{1u}=-\alpha(1-\alpha/2)J_{zz}/2$, $t_{2u}=t_{3u}=\alpha^{2}J_{zz}/4$. In addition, perturbative corrections give a momentum-independent energy shift, which arises from the disturbance induced by an excitation on the virtual transitions occurring in its proximity. This leads to a renormalization of the gap, which can be encoded in an energy shift $V_{u}=-5\alpha^{2}J_{zz}/8$. We note that the effective parameters $t_{1u}$, $t_{2u}$, $t_{3u}$, $V_{u}$ which we derived are equal, in a different notation, to those obtained in Ref. [8] for an equivalent problem of hard-core bosons.
@@
$\begin{split}\epsilon^{(2,{\rm LSWT})}_{u,\pm}(\mathbf{k})&=g_{z}\mu_{\rm B}H-\frac{S}{6}\alpha^{2}J_{zz}|f_{\mathbf{k}}|^{2}\\ &\qquad\pm S\alpha J_{zz}\left|f_{\mathbf{k}}-\frac{\alpha}{6}f^{*2}_{\mathbf{k}}\right|~{},\\ \epsilon^{(2,{\rm LSWT})}_{d}(\mathbf{k})&=6SJ_{zz}-g_{z}\mu_{\rm B}H-\frac{S}{3}\alpha^{2}J_{zz}|f_{\mathbf{k}}|^{2}~{}.\end{split}$ (5)
@@
Finally, we note that the expansion near the Ising limit offers the possibility to assess not only the quality of the LSWT approximation, but also the convergence of the semiclassical $1/S$ series. This can be controlled by studying the generalization of the problem to an arbitrary value of the spin $S$. To repeat the analysis, the $uud$ ground state configuration is promoted to a state in which two sublattices (A and B) are occupied by spins $+S$, whereas the remaining sublattice C is occupied by spins $-S$, and the $u$ and $d$ excitations, become the lowering or raising by $1$ of the spins on the A, B or on the C sublattices. At order $\alpha^{2}$, the effective hopping problem on the honeycomb lattice which governs the $u$ excitations has the same hopping amplitudes as in the $S=1/2$ case, but with a nontrivial dependence on the spin $S$:
@@
Data availability—The data that support the findings of this article are available [26].
@@
$\begin{split}\epsilon^{(2)}_{u,\pm}(\mathbf{k})&=g_{z}\mu_{\rm B}H+V_{u}+t_{1u}\Big{[}-\gamma(|f_{\mathbf{k}}|^{2}-3)\\ &\pm\left|(1-2\gamma)f_{\mathbf{k}}+\gamma f^{*2}_{\mathbf{k}}\right|\Big{]}~{},\end{split}$ (15)
@@
After applying the $\pi$ rotation $S_{i}^{x}\to S_{i}^{x}$, $S_{i}^{y}\to-S^{y}_{i}$, $S^{z}_{i}\to-S_{i}^{z}$ to the spins residing on the C sublattice, introducing the Holstein-Primakoff representations
@@
11footnotetext: See the Supplemental Materials of Ref. [22].22footnotetext: A. Mauri and F. Mila, ’Spin-wave expansion and magnon dispersion in the 1/3 plateau of the triangular XXZ antiferromagnet’. Zenodo, 2025. doi: https://doi.org/10.5281/zenodo.15270189